# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v2

@@ III-B Scene representation
In our MDP, a state consists of a large and varying number of agents. To aid policy learning with such a complicated state space, we develop a structured representation of states that will be fed into the structured policy network. To reduce sim2real gaps caused by raw sensor readings, our scene representation leverages processed information from perception, maps, and robot localization, which are relatively robust to domain shifts.
In Fig. 3, at each timestep $t$, we split a scene into a human representation ${h}^{t}_{1},...,{h}^{t}_{n}$, and an obstacle representation ${o}^{t}$. In human representation (Fig. 3(b)), the position and velocity of each human are detected using off-the-shelf human detectors [53, 54, 55]. Tracking each human is not required, further simplifying perception requirements. By representing each human as a low-dimentional state vector, we abstract away detailed information such as gaits and appearance, which are difficult to simulate accurately [38, 56].
Figure 3: A split representation of a crowded and constrained navigation scenario. In a dynamic scene, (b) human states are detected by perception modules. To represent static obstacles, we preprocess the original map (a) by combining and smoothing nearby objects to produce a clean contour. From this processed map and the robot’s position, we generate an artificial point cloud (c) that simulates obstacle perception. This representation allows the robot to learn robust navigation policies with smaller sim2real gaps in a low-fidelity simulator.
We use a 2D point cloud as the obstacle representation, which offers the following two variants. The first variant uses a real-time 2D point cloud obtained from onboard sensors. The second variant constructs an “artificial” point cloud which is computed from a pre-built map and the robot’s localization within that map. In this case, the point cloud is generated by raycasting from the robot’s estimated pose on the map, representing only static obstacles and unaffected by dynamic elements such as humans. The first variant reflects the robot’s instantaneous perception of its surroundings but requires a high-fidelity simulation of LiDAR data. The second variant is more robust to occlusions, imperfect human models, and sensor noise, but assumes the availability of an accurate map and reliable localization. In this work, we adopt the artificial point cloud variant to improve sim-to-real transferability.
@@ III-C Reward function
In summary, the reward function of our MDP is the sum of the above three parts:
@@
Intuitively, the robot gets a high reward when it approaches the goal with a high speed and a short and smooth path, while maintaining a safe distance from dynamic and static obstacles.
@@ IV-A Heterogeneous Spatio-Temporal Graph
The subtle yet highly dynamic interactions among agents and entities are important factors that makes crowd navigation difficult. To model these interactions in a structured manner, we formulate the navigation scenario as a heterogeneous st-graph. In Fig. 4a, at each timestep $t$, our heterogeneous st-graph $\mathcal{G}^{t}=(\mathcal{V}^{t},\mathcal{E}^{t})$ consists of a set of nodes $\mathcal{V}^{t}$ and a set of edges $\mathcal{E}^{t}$. The nodes include the detected humans $h_{1}^{t},...,h_{n}^{t}$ and the robot $w^{t}$. In addition, an obstacle node $o^{t}$ represents the point cloud of all obstacles as a whole. At each timestep $t$, the spatial edges that connect different nodes denote the spatial interactions among nodes. Different spatial interactions have different effects on robot decision-making. Specifically, since we have control of the robot but not the humans, RH interactions have direct effects while HH interactions have indirect effects on the robot actions. As an example of indirect effects, if human A aggressively forces human B to turn toward the robot’s front, the robot has to respond as a result of the interaction between A and B. Additionally, since the agents are dynamic but the obstacles are static, interactions among agents are mutual while the influence of static obstacles on agents is one-way. Thus, we categorize the spatial edges into three types: HH edges (blue in Fig. 4), obstacle-agent (OA) edges (orange), and RH edges (red). The three types of edges allow us to factorize the spatial interactions into HH, OA, and RH functions. Each function is parameterized by a neural network that has learnable parameters. Compared to previous works that ignore some edges [3, 8, 10], our method allows the robot to reason about all observed spatial interactions that exist in constrained and crowded environments.
@@
To reduce the number of parameters, the same type of edges in Fig. 4(a) share the same function parameters. This parameter sharing is important for the scalability of our graph because the number of parameters is kept constant when the number of human changes [57].
@@ IV-B1 Attention among agents
The attention modules assign weights to all edges that connect to a robot or a human node, allowing the node to pay attention to important edges or interactions. The two attention networks are similar to the scaled dot-product attention with a padding mask [44], which computes the attention score using a query $Q$ and a key $K$, and applies the normalized score to a value $V$, which results in a weighted value $v$.
@@
In addition, following Eq. 4, a mask $M^{t}\in\mathbb{R}^{n\times n}$ indicates the visibility of each human to the robot at current time $t$ and is obtained from the perception system of the robot. Assume the $n$-th human is not visible at time $t$. Then $M^{t}$ is a matrix filled with zeros, except that every entry in $n$-th column is $\infty$- ∞. The numerator of Eq. 4 can be express as
@@
where $c_{1},...,c_{n}$ are constants that reflect the effect of the scaling factor $d$ and the softmax function. Thus, the value of the $n$-th missing human $v_{n}$ is eliminated by the mask $M^{t}$ and thus does not affect the resulting weighted human embedding $v_{HH}^{t}$. The mask that indicates the visibility of each human prevents attention to undetected humans, which is common in crowd navigation due to the partial observability caused by the limited robot sensor range, occlusions, imperfect human detectors, etc [33]. Additionally, the mask provides unbiased gradients to the networks, which stabilizes and accelerates the training [58, 10].
@@ IV-B2 Incoporating obstacle and temporal information
Finally, the robot and obstacle embeddings are concatenated with the twice weighted human features $v_{RH}^{t}$ and fed into the GRU [^1] From experiments, we find that adding a third obstacle-agent attention network, where the obstacle embedding is the key and the robot and human embeddings are the query and value, does not improve the performance. Thus, we keep the simple concatenation design to incorporate obstacle information.:
@@ IV-B3 Training
We train the entire network with Proximal Policy Optimization (PPO) [59] in a simulator as shown in Fig. 2(a) (see Sec. V-A for details of the simulator). At each timestep $t$, the simulator provides all state information that constitutes $s^{t}$, which is fed to the HEIGHT network. The network outputs the estimated value of the state $V(s^{t})$ and the logarithmic probabilities of the robot’s action $\pi(a^{t}|s^{t})$, both of which are used to compute the PPO loss and then update the parameters in the network. In training, the robot action is sampled from the action distribution $\pi(a^{t}|s^{t})$. In testing, the robot takes the action with the highest probability $a^{t}$. The robot action $a^{t}$ is fed into the simulator to compute the next state $s^{t+1}$, and then the loop continues.
Without any supervised learning, our method is not limited by the performance of expert demonstrations [4, 3, 51]. However, to improve the low training data efficiency, an inherent problem of RL, HEIGHT can also be trained with a combination of imitation learning and RL.
@@ V Simulation Experiments
Method Scene Representation Navigation algorithm
DWA [1] Groups of small circles Treats humans as obstacles, no interaction modeling
$\textrm{A}^{*}$+ CNN [7] LiDAR point cloud A* + RL policy with no interaction modeling
DS-RNN [8] Groups of small circles RL policy with only RH attention and GRU
DRL-VO [30] Occupancy map + LiDAR point cloud Pure pursuit algorithm + RL policy with no interaction modeling
HomoGAT [45] Obstacle point cloud + detected human circles RL policy derived from st-graph with shared attention weights for all edges
Ours Obstacle point cloud + detected human circles RL policy derived from st-graph with separate functions for all edges
@@
V-A Simulation environment
We conduct simulation experiments in random environment in Fig. 5(a) developed with PyBullet [60]. The robot, humans, and static obstacles are in a 12 m×$\times$12 m arena. In each episode, rectangular obstacles are initialized with random shapes and random poses. The width and the length of each static obstacle are sampled from $\mathcal{N}(1,0.6^{2})$ and are limited in $[0.1,5]$ meters. The initial positions of the humans and the robot are also randomized. The starting and goal positions of the robot are randomly sampled inside the arena. The distance between the robot’s starting and the goal positions is between 5 m and 6 m. Some humans are dynamic and some are static. The goals of dynamic humans are set on the opposite side of their initial positions to create circle-crossing scenarios. In training, the number of humans varies from 5 to 9 and the number of static rectangular obstacles varies from 8 to 12. Among all humans, 0-2 of them are static and the rest are dynamic. In testing, the number of humans and obstacles are shown in the first column of Table II. The heights of humans and obstacles are all above the height of the robot LiDAR to ensure detectability.
To simulate a continuous human flow similar to [8, 10, 30], dynamic humans will move to new random goals immediately after they arrive at their goal positions or they get stuck in front of narrow passageways for more than 10 timesteps. All dynamic humans are controlled by ORCA [15]. 80% of dynamic humans do not react to the robot and 20% of humans react to the robot. This mixed setting prevents our network from learning an extremely aggressive policy in which the robot forces all humans to yield while achieving a high reward. Meanwhile, the simulation maintains enough reactive humans to resemble the real crowd behaviors. We use holonomic kinematics for humans. The preferred speed of humans ranges from 0.4 m/s to 0.6 m/s to accommodate the speed of the robot. We assume that humans can achieve the desired velocities immediately, and they will keep moving with these velocities for the next $\Delta t$ seconds. The simulation and control frequency $\Delta t$ is 0.1 s. The maximum length of an episode is 491 timesteps or 49.1 s.
Figure 5: Simulation and real-world environments. (a) In our simulation benchmark, the colored cylinders denote humans and gray objects denote obstacles. Each episode is randomized (Sec. V-A). (b) and (c) are everyday indoor environments for sim2real. Before real-world testing in (b) and (c), the robot policy is trained in (d) and (e) respectively (Sec. VI-A).
V-B Experiment setup
We now introduce the baselines and ablation models, training procedure, and evaluation metrics.
@@ V-B2 Ablations
[…]
Overall w/ Humans w/ Obstacles
Training distribution 5-9 humans 8-12 obstacles DWA [1] 0.16 0.68 0.63 0.05 0.15 28.45 11.52
$\textrm{A}^{*}$+ CNN [7] 0.64 0.29 0.28 0.01 0.07 25.72 12.30
DS-RNN [8] 0.80 0.20 0.11 0.09 0.00 19.90 10.78
DRL-VO [30] 0.59 0.41 0.34 0.07 0.00 21.45 10.36
HomoGAT [45] 0.86 0.14 0.13 0.01 0.00 18.66 10.36
HEIGHT (ours) 0.88 0.12 0.09 0.03 0.00 18.31 10.34
Less crowded 0-4 humans 8-12 obstacles DWA [1] 0.29 0.37 0.28 0.09 0.34 25.74 14.42
$\textrm{A}^{*}$+ CNN [7] 0.84 0.12 0.11 0.01 0.04 22.64 11.73
DS-RNN [8] 0.72 0.28 0.07 0.21 0.00 17.21 9.56
DRL-VO [30] 0.82 0.17 0.12 0.05 0.01 19.14 10.53
HomoGAT [45] 0.95 0.05 0.03 0.02 0.00 16.64 10.17
HEIGHT (ours) 0.97 0.03 0.02 0.01 0.00 16.32 10.04
More crowded 10-14 humans 8-12 obstacles DWA [1] 0.11 0.78 0.74 0.04 0.11 29.60 10.37
$\textrm{A}^{*}$+ CNN [7] 0.47 0.42 0.39 0.03 0.11 27.33 12.47
DS-RNN [8] 0.76 0.22 0.16 0.06 0.01 23.08 11.67
DRL-VO [30] 0.50 0.49 0.41 0.09 0.01 22.72 10.26
HomoGAT [45] 0.72 0.28 0.23 0.05 0.00 20.46 10.57
HEIGHT (ours) 0.78 0.22 0.19 0.03 0.00 19.69 10.39
Less constrained 5-9 humans 3-7 obstacles DWA [1] 0.17 0.57 0.55 0.02 0.26 28.50 14.21
$\textrm{A}^{*}$+ CNN [7] 0.66 0.29 0.28 0.01 0.05 23.63 11.98
DS-RNN [8] 0.66 0.34 0.20 0.14 0.00 18.01 9.76
DRL-VO [30] 0.70 0.30 0.26 0.04 0.00 21.21 11.15
HomoGAT [45] 0.88 0.12 0.10 0.02 0.00 17.66 10.37
HEIGHT (ours) 0.90 0.10 0.09 0.01 0.00 17.20 10.23
More constrained 5-9 humans 13-17 obstacles DWA [1] 0.14 0.66 0.49 0.17 0.20 30.47 11.35
$\textrm{A}^{*}$+ CNN [7] 0.48 0.29 0.23 0.06 0.23 27.28 13.11
DS-RNN [8] 0.71 0.23 0.09 0.14 0.06 23.45 11.58
DRL-VO [30] 0.55 0.40 0.23 0.07 0.05 21.87 10.22
HomoGAT [45] 0.80 0.20 0.11 0.09 0.00 19.20 10.51
HEIGHT (ours) 0.84 0.15 0.07 0.08 0.01 18.79 10.65
@@ V-B3 Training
For fair comparision, we train all RL methods, including all baseines and ablations except DWA, with the same number of timesteps and the same PPO hyperparameters. Please refer to Appendix-B for training details.

### v3

@@ III-B Scene representation
To aid policy learning with the complicated state space involving multiple agents and entities, we develop a structured representation that splits human and obstacle states to account for their different effects on robot decision making. In Fig. 3, at each timestep $t$, we split a scene into a human representation ${h}^{t}_{1},...,{h}^{t}_{n}$, and an obstacle representation ${o}^{t}$. Each human has its own state since it can move independently from one another, whereas all obstacles are mixed together since they are all static. In human representation (Fig. 3(b)), the position and velocity of each human are detected using off-the-shelf human detectors [Jia2020DRSPAAM, yolov3, deepsort]. Tracking each human is not required, further simplifying perception requirements. By representing each human as a low-dimentional state vector, we abstract away detailed information such as gaits and appearance, which are difficult to simulate accurately [tsoi2022sean2, mavrogiannis2023core].
Figure 3: A split representation of a crowded and constrained navigation scenario. In (b), individual human states are represented by circles. To represent static obstacles, we preprocess and smooth the original map (a). From this processed map and the robot’s localization, we generate an artificial point cloud (c) that simulates obstacle perception.
We use a 2D point cloud as the obstacle representation, which offers the following two variants. The first variant uses a real-time 2D point cloud obtained from onboard sensors. The second variant constructs an “artificial” point cloud which is computed from a pre-built map and the robot’s localization in the map. In this case, the point cloud is generated by raycasting from the robot’s estimated pose on the map, representing only static obstacles and unaffected by humans. The first variant reflects the robot’s instantaneous perception of its surroundings but requires a high-fidelity simulation of LiDAR data. The second variant is more robust to occlusions, imperfect human models, and sensor noise, but assumes the availability of a map and reliable localization. In this work, since we have access to reliable robot localization and a low-fidelity simulator with inaccurate obstacle shapes and LiDAR point clouds, we opt for the second choice to reduce sim2real gap and improve the poilcy’s real-world performance.
@@ III-C Reward function
The robot’s reward function is the sum of the three parts:
@@
Intuitively, the robot gets a high reward when it approaches the goal with a high speed and a short and smooth path, while maintaining a safe distance from dynamic and static obstacles. The coefficient of each reward term indicates the priority of each subtask (e.x. reaching the goal and avoiding collisions are more important than others) and can be tuned if the desired priority changes.
@@ IV-A Heterogeneous Spatio-Temporal Graph
The subtle yet highly dynamic interactions among agents and entities are important factors that makes crowd navigation difficult. To model these interactions in a structured manner, we formulate the navigation scenario as a heterogeneous st-graph. In Fig. 4a, at each timestep $t$, our heterogeneous st-graph $\mathcal{G}^{t}=(\mathcal{V}^{t},\mathcal{E}^{t})$ consists of a set of nodes $\mathcal{V}^{t}$ and a set of edges $\mathcal{E}^{t}$. The nodes include the detected humans $h_{1}^{t},...,h_{n}^{t}$ and the robot $w^{t}$. In addition, an obstacle node $o^{t}$ represents the point cloud of all obstacles as a whole. At each timestep $t$, the spatial edges that connect different nodes denote the spatial interactions among nodes. Different spatial interactions have different effects on robot decision-making. Specifically, since we have control of the robot but not the humans, RH interactions have direct effects while HH interactions have indirect effects on the robot actions. As an example of indirect effects, if human A aggressively forces human B to turn toward the robot’s front, the robot has to respond as a result of the interaction between A and B. Additionally, since the agents are dynamic but the obstacles are static, interactions among agents are mutual while the influence of static obstacles on agents is one-way. Thus, we categorize the spatial edges into three types: HH edges (blue in Fig. 4), obstacle-agent (OA) edges (orange), and RH edges (red). The three types of edges allow us to factorize the spatial interactions into HH, OA, and RH functions. Each function is parameterized by a neural network that has learnable parameters. Compared to previous works that ignore some edges [chen2019crowd, liu2020decentralized, liu2023intention], our method allows the robot to reason about all observed spatial interactions that exist in constrained and crowded environments.
@@
To reduce the number of parameters, the same type of edges in Fig. 4(a) share the same function parameters. This parameter sharing is important for the scalability of our graph because the number of parameters is kept constant when the number of human changes [jain2016structural].
@@ IV-B1 Attention among agents
The attention modules assign weights to all edges that connect to a robot or a human node, allowing the node to pay attention to important edges or interactions. The two attention networks are similar to the scaled dot-product attention with a padding mask [vaswani2017attention], which computes the attention score using a query $Q$ and a key $K$, and applies the normalized score to a value $V$, which results in a weighted value $v$.
@@
In addition, following Eq. 4, a mask $M^{t}\in\mathbb{R}^{n\times n}$ indicates the visibility of each human to the robot at current time $t$ and masks out humans that are undetectable by the robot sensors. Assume the $n$-th human is not visible at time $t$. Then $M^{t}$ is a matrix filled with zeros, except that every entry in $n$-th column is -$\infty$. The numerator of Eq. 4 can be express as
@@
where $c_{1},...,c_{n}$ are constants that reflect the effect of the scaling factor $d$ and the softmax function. Thus, the value of the $n$-th missing human $v_{n}$ is eliminated by the mask $M^{t}$ and thus does not affect the resulting weighted human embedding $v_{HH}^{t}$. The mask that indicates the visibility of each human prevents attention to undetected humans, which is common in crowd navigation due to the partial observability caused by the limited robot sensor range, occlusions, imperfect human detectors, etc [mun2023occlusion]. Additionally, the mask provides unbiased gradients to the networks, which stabilizes and accelerates the training [ma2020reinforcement, liu2023intention].
@@ IV-B2 Incoporating obstacle and temporal information
Finally, the robot and obstacle embeddings are concatenated with the twice weighted human features $v_{RH}^{t}$ and fed into the GRU [^1] From experiments, we find that adding a third obstacle-agent attention network, where the obstacle embedding is the key and the robot and human embeddings are the query and value, does not improve the performance. The reason is that the agents only interact with a small subset of obstacles instead of the whole point cloud at each timestep, eliminating the need for modeling these interactions. Thus, we keep the simple concatenation design to incorporate obstacle information.:
@@ IV-B3 Training
We train the entire network with Proximal Policy Optimization (PPO) [schulman2017proximal] in a simulator as shown in Fig. 2(a) (see Sec. V-A for details of the simulator). At each timestep $t$, the simulator provides all state information that constitutes $s^{t}$, which is fed to the HEIGHT network. The network outputs the estimated value of the state $V(s^{t})$ and the logarithmic probabilities of the robot’s action $\pi(a^{t}|s^{t})$, both of which are used to compute the PPO loss and then update the parameters in the network. In training, the robot action is sampled from the action distribution $\pi(a^{t}|s^{t})$. In testing, the robot takes the action with the highest probability $a^{t}$. The robot action $a^{t}$ is fed into the simulator to compute the next state $s^{t+1}$, and then the loop continues.
Without any supervised learning, our method is not limited by the performance of expert demonstrations [chen2017decentralized, chen2019crowd, liu2024sample]. However, to improve the low training data efficiency, an inherent problem of RL, HEIGHT can also be trained with a combination of imitation learning and RL.
@@ V Simulation Experiments
V-A Simulation benchmark
We conduct simulation experiments in a PyBullet simulator [coumans2019]. The robot, humans, and static obstacles are in a $\times12 m$ arena. In each episode, rectangular obstacles are initialized with random shapes and random poses. The width and the length of each static obstacle are sampled from $\mathcal{N}(1,0.6^{2})$ and are limited in $[0.1,5]$ meters. The initial positions of the humans and the robot are also randomized. The starting and goal positions of the robot are randomly sampled inside the arena. The distance between the robot’s starting and the goal positions is between 5 m and 6 m. Some humans are dynamic and some are static. The goals of dynamic humans are set on the opposite side of their initial positions to create circle-crossing scenarios. In training, the number of humans varies from 5 to 9 and the number of static rectangular obstacles varies from 8 to 12. Among all humans, 0-2 of them are static and the rest are dynamic. In testing, the number of humans and obstacles are shown in the first column of Table II. The heights of humans and obstacles are all above the height of the robot LiDAR to ensure detectability.
To simulate a continuous human flow similar to [liu2020decentralized, liu2023intention, xie2023drlvo], dynamic humans will move to new random goals immediately after they arrive at their goal positions or they get stuck in front of narrow passageways for more than 10 timesteps. All dynamic humans are controlled by ORCA [van2011reciprocal]. 80% of dynamic humans do not react to the robot and 20% of humans react to the robot. This mixed setting prevents our network from learning an extremely aggressive policy in which the robot forces all humans to yield while achieving a high reward. Meanwhile, the simulation maintains enough reactive humans to resemble the real crowd behaviors. We use holonomic kinematics for humans. The preferred speed of humans ranges from 0.4 m/s to 0.6 m/s to accommodate the speed of the robot. We assume that humans can achieve the desired velocities immediately, and they will keep moving with these velocities for the next $\Delta t$ seconds. The simulation and control frequency $\Delta t$ is 0.1 s. The maximum length of an episode is 491 timesteps or 49.1 s.
V-B Setup
@@ V-B1 Baselines
Method Scene Representation (humans and static obstacles) Navigation algorithm
DWA [fox1997dynamic] Circles for both Treats humans as obstacles, no interaction modeling
ORCA [van2011reciprocal] Circles for humans + Line segments for obstacles Optimization program with velocity obstacles to model all interactions
$\textrm{A}^{*}$+ CNN [perez2021robot] LiDAR point cloud for both A* + RL policy with no interaction modeling
DS-RNN [liu2020decentralized] Circles for both RL policy with only RH attention and GRU
DRL-VO [xie2023drlvo] Occupancy map for humans + Point cloud for both Pure pursuit algorithm + RL policy with no interaction modeling
HomoGAT [velickovic2018graph] Circles for humans + Point cloud for obstacles RL policy from st-graph with shared attention weights for all interactions
Ours Circles for humans + Point cloud for obstacles RL policy from st-graph with separate functions for all interactions
@@ V-B3 Training
For fair comparision, we train all RL methods, including all baseines and ablations except DWA, with the same number of timesteps and the same PPO hyperparameters. Please refer to Appendix-B-1) for training details.
@@ V-B4 Evaluation
[…]
Overall w/ Humans w/ Obstacles
Training distribution 5-9 humans 8-12 obstacles DWA [fox1997dynamic] 0.16 0.68 0.63 0.05 0.15 28.45 11.52
ORCA [van2011reciprocal] 0.34 0.62 0.31 0.32 0.04 21.51 9.46
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.64 0.29 0.28 0.01 0.07 25.72 12.30
DS-RNN [liu2020decentralized] 0.80 0.20 0.11 0.09 0.00 19.90 10.78
DRL-VO [xie2023drlvo] 0.59 0.41 0.34 0.07 0.00 21.45 10.36
HomoGAT [velickovic2018graph] 0.86 0.14 0.13 0.01 0.00 18.66 10.36
HEIGHT (ours) 0.88 0.12 0.09 0.03 0.00 18.31 10.34
Less crowded 0-4 humans 8-12 obstacles DWA [fox1997dynamic] 0.29 0.37 0.28 0.09 0.34 25.74 14.42
ORCA [van2011reciprocal] 0.42 0.54 0.11 0.42 0.04 19.77 9.69
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.84 0.12 0.11 0.01 0.04 22.64 11.73
DS-RNN [liu2020decentralized] 0.72 0.28 0.07 0.21 0.00 17.21 9.56
DRL-VO [xie2023drlvo] 0.82 0.17 0.12 0.05 0.01 19.14 10.53
HomoGAT [velickovic2018graph] 0.95 0.05 0.03 0.02 0.00 16.64 10.17
HEIGHT (ours) 0.97 0.03 0.02 0.01 0.00 16.32 10.04
More crowded 10-14 humans 8-12 obstacles DWA [fox1997dynamic] 0.11 0.78 0.74 0.04 0.11 29.60 10.37
ORCA [van2011reciprocal] 0.26 0.71 0.41 0.30 0.03 23.33 9.21
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.47 0.42 0.39 0.03 0.11 27.33 12.47
DS-RNN [liu2020decentralized] 0.76 0.22 0.16 0.06 0.01 23.08 11.67
DRL-VO [xie2023drlvo] 0.50 0.49 0.41 0.09 0.01 22.72 10.26
HomoGAT [velickovic2018graph] 0.72 0.28 0.23 0.05 0.00 20.46 10.57
HEIGHT (ours) 0.78 0.22 0.19 0.03 0.00 19.69 10.39
Less constrained 5-9 humans 3-7 obstacles DWA [fox1997dynamic] 0.17 0.57 0.55 0.02 0.26 28.50 14.21
ORCA [van2011reciprocal] 0.42 0.54 0.30 0.24 0.03 21.89 10.59
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.66 0.29 0.28 0.01 0.05 23.63 11.98
DS-RNN [liu2020decentralized] 0.66 0.34 0.20 0.14 0.00 18.01 9.76
DRL-VO [xie2023drlvo] 0.70 0.30 0.26 0.04 0.00 21.21 11.15
HomoGAT [velickovic2018graph] 0.88 0.12 0.10 0.02 0.00 17.66 10.37
HEIGHT (ours) 0.90 0.10 0.09 0.01 0.00 17.20 10.23
More constrained 5-9 humans 13-17 obstacles DWA [fox1997dynamic] 0.14 0.66 0.49 0.17 0.20 30.47 11.35
ORCA [van2011reciprocal] 0.30 0.58 0.24 0.34 0.11 22.40 9.04
$\textrm{A}^{*}$+ CNN [perez2021robot] 0.48 0.29 0.23 0.06 0.23 27.28 13.11
DS-RNN [liu2020decentralized] 0.71 0.23 0.09 0.14 0.06 23.45 11.58
DRL-VO [xie2023drlvo] 0.55 0.40 0.23 0.07 0.05 21.87 10.22
HomoGAT [velickovic2018graph] 0.80 0.20 0.11 0.09 0.00 19.20 10.51
HEIGHT (ours) 0.84 0.15 0.07 0.08 0.01 18.79 10.65